OpenStudy (anonymous):

Calc 3 help! Suppose that Pi + Qj is parallel to the tangent vector of a closed curve C. 2 questions about this! Please help!

OpenStudy (anonymous):

a) Show that Qi-Pj is perpendicular to the tangent vector. b) Show that \[\int\limits_{}^{}\int\limits_{D}^{}(\delta P/\delta x + \delta Q/\delta y)dxdy = 0\] where D is the region whose boundary is C

OpenStudy (turingtest):

to show that Qi-Pj is perpendicular to Pi+Qj just take the dot product

OpenStudy (anonymous):

so it should be 0, right? what about part b?

OpenStudy (turingtest):

still thinking on that part, you sure that's not supposed to be a - sign in the double integral?

OpenStudy (anonymous):

Yeah, I wish it was too... But its supposed to be a +

OpenStudy (turingtest):

\[\iint\limits_D\left(\frac{\partial P}{\partial x}+\frac{\partial Q}{\partial y}\right)dA=\oint\limits_CQdx-Pdy\]since we have established that Qi-Pj is perpendicular to every point on C, clearly this integral will be zero. I don't see a more rigorous way to prove it just yet.

OpenStudy (turingtest):

does that at least make sense though?

OpenStudy (turingtest):

still there?

OpenStudy (anonymous):

Yeah, I'm back. That makes sense a little bit actually. Thank you!